OpenStudy (anonymous):

A)14 is 30 percent of what number? B) 72 is what percent of 180?

OpenStudy (anonymous):

thx in b would u divide as well.

OpenStudy (anonymous):

Well, 140/3=46.67, not 4.6

OpenStudy (anonymous):

\[\frac{ 140 }{ 3 }=46\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

in B would u divide or X

OpenStudy (anonymous):

Use this general statement for (b)